OpenStudy (anonymous):

13. What is the inverse operations of subtraction? A. multiplication B. division C. addition D. subtraction 14. Choose ALL of the words below that mean "equal." A. is B. the same as C. of D. equal to E. increased F. identical to 15. Sam needs to make 156 cookies for a school bake sale. How many dozens does he need to make? (1 dozen = 12) A. 12 B. 168 C. 13 D. 144 16. What is the goal when solving an equation? A. use inverse operations B. isolate the variable C. add and subtract D. get the variable on the right

OpenStudy (anonymous):

17. Solve: x + 15 = 8 A. x = 7 B. x = 23 C. x = -23 D. x = -7 18. Solve: y - 12 = 34 A. y = 46 B. y = 22 C. y - 12 + 12 = 34 + 12 D. y = 34 19. Solve: 72 + c = 55 A. c = 127 B. c = 17 C. c = -17 D. none of the above